# Continuous-time q-learning in jump-diffusion models under Tsallis entropy

### v3

@@ 3 Continuous-time q-Function and Martingale Characterization under Tsallis Entropy
Suppose (A1), (A2) and (A3) hold. Given $(t,x)\in[0,T]\times\mathds{R}$ and $\pi,\pi^{\prime}\in\Pi_{t}$, if it holds that $F(t,x;\pi^{\prime},\pi)\geq F(t,x;\pi,\pi)$, then $J(t,x;\pi^{\prime})\geq J(t,x;\pi)$.
Moreover, in order to employ the q-learning method based on Theorem 3.3, the policy function $\pi^{\xi}$ should satisfy $\pi^{\chi}\in{\cal P}(U)$ and the consistency condition (3.6). Here, we relax these constraints and consider the following maximization problem, for $w_{1},w_{2}\geq 0$
@@
1:Initialize $\theta$$\zeta$ and $i=1$.
@@ 5.1 The optimal portfolio liquidation problem
Consider an optimal portfolio liquidation problem in which a large investor has access both to a classical exchange and to a dark pool with adverse selection. As in Kratz and Schöneborn (2014, 2015), the trading and price formation is described as the classical exchange as a linear price impact model. The trade execution can be enforced by selling aggressively, which however results in quadratic execution costs due to a stronger price impact. The order execution in the dark pool is modeled by a Poisson process ${N}=(N_{t})_{t\geq 0}$ with intensity parameter $\lambda>0$, where orders submitted to the dark pool are executed at the jump times of Poisson processes. The split of orders between the dark pool and exchange is thus driven by the trade-off between execution uncertainty and price impact costs. Next, we formulate the optimal portfolio liquidation problem in detail. Consider an investor who has to liquidate an asset position $x\in\mathds{R}$ within a finite trading horizon $[0,T]$. The investor can control her trading intensity $\xi=(\xi_{t})_{t\in[0,T]}$, and they can place orders $\eta=(\eta_{t})_{t\in[0,T]}$ in the dark pool. Given a trading strategy $u=(\xi_{t},\eta_{t})_{t\in[0,T]}$ (as r.c.l.l. $\mathbb{F}$-predictable processes) taking values on the policy space $U=\mathds{R}^{2}$, the asset holdings of the investor at time $t\in[0,T)$ is given by
@@
$\begin{cases}\displaystyle w(t,x):=\sup_{\pi\in\Pi}\mathbb{E}^{\mathbb{P}}\left[\int_{t}^{T}\int_{U}\left(-\kappa|u_{1}|^{2}-c|X_{s}^{u}|^{2}+\gamma l_{p}(\pi_{s}(u))\right)\pi_{s}(u)duds-g_{T}(X_{T}^{u})\Big{|}X_{t}=x\right],\\[11.99998pt] \displaystyle\leavevmode\nobreak\ {\rm s.t.}\leavevmode\nobreak\ X_{t}^{\pi}=x-\int_{0}^{t}\int_{\mathds{R}^{2}}u_{1}\pi_{s}(du)ds-\int_{0}^{t}\int_{\mathds{R}^{2}}u_{2}\mathcal{N}(ds,du),\quad\forall t\in[0,T].\end{cases}$ (5.5)
@@
$\begin{cases}\displaystyle V^{(\ell)}(t,x):=\sup_{\pi\in\Pi}J^{(\ell)}(t,x;\pi)\\[8.00003pt] \displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ =\sup_{\pi\in\Pi}\mathbb{E}^{\mathbb{P}}\left[\int_{t}^{T}\int_{U}\left(-\kappa|u_{1}|^{2}-c|X_{s}^{u}|^{2}+\gamma l_{p}(\pi_{s}(u))\right)\pi_{s}(u)duds-\ell X_{T}^{2}\Big{|}X_{t}=x\right],\\[11.99998pt] \displaystyle\leavevmode\nobreak\ {\rm s.t.}\leavevmode\nobreak\ X_{t}^{\pi}=x-\int_{0}^{t}\int_{\mathds{R}^{2}}u_{1}\pi_{s}(du)ds-\int_{0}^{t}\int_{\mathds{R}^{2}}u_{2}\mathcal{N}(ds,du),\quad\forall t\in[0,T].\displaystyle\end{cases}$ (5.6)
@@
$\widehat{\pi}^{(\ell)}(u|t,x)=\begin{cases}\displaystyle\left(\frac{p-1}{\gamma p}\right)^{\frac{1}{p-1}}\left(\psi(t,x)-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}+\frac{\gamma}{p-1}\right)_{+}^{\frac{1}{p-1}},\leavevmode\nobreak\ p>1,\\ \\ \displaystyle\frac{\exp\left(-u_{1}V_{x}^{\ell}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}-\gamma\right)}{\int_{\mathds{R}^{2}}\exp\left(-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}-\gamma\right)du},\leavevmode\nobreak\ p=1.\end{cases}$
@@
$\widehat{\pi}^{(\ell)}(u|t,x)=\begin{cases}\displaystyle\left(\frac{p-1}{\gamma p}\right)^{\frac{1}{p-1}}\left(\psi(t,x)-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}+\frac{\gamma}{p-1}\right)_{+}^{\frac{1}{p-1}},\leavevmode\nobreak\ p>1,\\[6.80011pt] \displaystyle\frac{\exp\left(-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}-\gamma\right)}{\int_{\mathds{R}^{2}}\exp\left(-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}-\gamma\right)du},\leavevmode\nobreak\ p=1\end{cases}$
@@
$=\begin{cases}-\frac{\alpha^{(\ell)}(t)x}{2\kappa},\leavevmode\nobreak\ m=1,\\ \ \\ \frac{\tilde{\psi}(t)(p-1)}{2\kappa(2p-1)}+\frac{(\alpha^{(\ell)}(t)x)^{2}}{4\kappa^{2}},\leavevmode\nobreak\ m=2,\end{cases}$
@@
$\leavevmode\nobreak\ \int_{\mathds{R}^{2}}\frac{1}{p-1}\left(\pi^{(\ell)}(u|t,x)-\pi^{(\ell)}(u|t,x)^{p}\right)du=\frac{1}{p-1}-\int_{\mathds{R}^{2}}\frac{\pi^{(\ell)}(u|t,x)^{p}}{p-1}du$
@@
$\begin{cases}\displaystyle{\alpha^{(\ell)}_{t}(t)}=-\frac{(\alpha^{(\ell)}(t))^{2}}{2\kappa}+{\lambda}\alpha^{(\ell)}(t)+2c,\leavevmode\nobreak\ \leavevmode\nobreak\ \alpha^{(\ell)}(T)=-\ell,\\[8.00003pt] \displaystyle\beta^{(\ell)}_{t}(t)=\frac{\tilde{\psi}(t)(p-1)}{2p-1}-\gamma\left(\frac{1}{p-1}-\frac{\tilde{\psi}(t)}{(2p-1)\gamma}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ \beta^{(\ell)}(T)=0.\end{cases}$
@@
For the case with $p=1$, the optimal policy $\widehat{\pi}$ given by (5.8) is a two-dimensional Gaussian distribution; while for $p>1$, the optimal policy becomes a two-dimensional q-Gaussian distribution with a compact support set, see Figure 1 for the illustration. In fact, for $p>1$ and $(t,x)\in[0,T]\times\mathds{R}_{+}$, we have
@@
Figure 1: (a) The optimal policy $(u_{1},u_{2})\to\hat{\pi}(u_{1},u_{2})$ with $p=1$. (b): The optimal policy $(u_{1},u_{2})\to\hat{\pi}(u_{1},u_{2})$ with $p=2$. (c): The optimal policy $(u_{1},u_{2})\to\hat{\pi}(u_{1},u_{2})$ with $p=3$. The model parameters are set to be $\lambda=1,\leavevmode\nobreak\ \kappa=1,\leavevmode\nobreak\ c=1,\leavevmode\nobreak\ \gamma=1,\leavevmode\nobreak\ t=1,\leavevmode\nobreak\ T=2,\leavevmode\nobreak\ x=5$.
@@
which can also be observed from Figure 2 and Figure 3.
Figure 2: (a): The compact support set of $u_{1}$. (b) The compact support set of $u_{2}$. The model parameters are set to be $\lambda=1,\leavevmode\nobreak\ \kappa=1,\leavevmode\nobreak\ c=1,\leavevmode\nobreak\ \gamma=0.001,\leavevmode\nobreak\ t=1,\leavevmode\nobreak\ T=2,\leavevmode\nobreak\ x=5$.
Figure 3: (a): The compact support set of $u_{1}$. (b) The compact support set of $u_{2}$. The model parameters are set to be $\lambda=1,\leavevmode\nobreak\ \kappa=1,\leavevmode\nobreak\ c=1,\leavevmode\nobreak\ \gamma=0.1,\leavevmode\nobreak\ t=1,\leavevmode\nobreak\ T=2,\leavevmode\nobreak\ x=5$.
@@
which yields the convergence of the optimal trading policy to a constant strategy that $(\xi_{t},\eta_{t})\xrightarrow[\gamma\to 0]{L^{2}}\left(\frac{\alpha^{*}(t)x}{2\kappa},x\right)$ for all $(t,x)\in[0,T]\times\mathds{R}$.
@@
Simulator: In what follows, we apply Algorithm 1 with the above parameterized value function and q-function. To generate sample trajectories, we first apply the acceptance-rejection sampling method (Flury 1990) to generate the control pair $(u^{1}_{t},u^{2}_{t})$ from the q-Gaussian distribution with density function given by (5.12) at time $t$. Then, the control pair $(u^{1}_{t},u^{2}_{t})$ is used to the following simulator
@@
$\alpha_{\theta_{1}}(k)$ $=\begin{cases}\displaystyle 0.01,\leavevmode\nobreak\ \text{if }1\leq k\leq 2500,\\ \displaystyle\frac{0.001}{{\rm linspace}_{(1,20,N)}(k)},\leavevmode\nobreak\ \text{if }2500<k\leq N,\end{cases}\leavevmode\nobreak\ \alpha_{\theta_{2}}(k)=\begin{cases}\displaystyle 0.005,\leavevmode\nobreak\ \text{if }1\leq k\leq 4000,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,100,N)}(k)},\leavevmode\nobreak\ \text{if }4000<k\leq N,\end{cases}$
$\alpha_{\theta_{3}}(k)$ $=\begin{cases}\displaystyle 0.01,\leavevmode\nobreak\ \text{if }1\leq k\leq 4000,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,20,N)}(k)},\leavevmode\nobreak\ \text{if }4000<k\leq N,\end{cases}\leavevmode\nobreak\ \alpha_{\theta_{4}}(k)=\begin{cases}\displaystyle 0.03,\leavevmode\nobreak\ \text{if }1\leq k\leq 3000,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,20,N)}(k)},\leavevmode\nobreak\ \text{if }3000<k\leq N,\end{cases}$
$\alpha_{\theta_{5}}(k)$ $=\begin{cases}\displaystyle 0.05,\leavevmode\nobreak\ \text{if }1\leq k\leq 3000,\\ \displaystyle\frac{0.0005}{{\rm linspace}_{(1,20,N)}(k)},\leavevmode\nobreak\ \text{if }3000<k\leq N,\end{cases}\leavevmode\nobreak\ \alpha_{\zeta_{1}}(k)=\begin{cases}\displaystyle 0.03,\leavevmode\nobreak\ \text{if }1\leq k\leq 3500,\\ \displaystyle\frac{0.00135}{{\rm linspace}_{(1,10,N)}(k)},\leavevmode\nobreak\ \text{if }3500<k\leq N,\end{cases}$
$\alpha_{\zeta_{2}}(k)$ $=\begin{cases}\displaystyle 0.1,\leavevmode\nobreak\ \text{if }1\leq k\leq 3500,\\ \displaystyle\frac{0.0002}{{\rm linspace}_{(1,500,N)}(k)},\leavevmode\nobreak\ \text{if }3500<k\leq N,\end{cases}\leavevmode\nobreak\ \alpha_{\zeta_{3}}(k)=\begin{cases}\displaystyle 0.1,\leavevmode\nobreak\ \text{if }1\leq k\leq 2000,\\ \displaystyle 0.002,\leavevmode\nobreak\ \text{if }2000<k\geq 5000,\\ \displaystyle\frac{0.0005}{{\rm linspace}_{(1,20,N)}(k)},\leavevmode\nobreak\ \text{if }k\geq 5000,\end{cases}$
$\alpha_{\zeta_{4}}(k)$ $=\begin{cases}\displaystyle 0.005,\leavevmode\nobreak\ \text{if }1\leq k\leq 7000,\\ \displaystyle\frac{0.001}{{\rm linspace}_{(1,100,N)}(k)},\leavevmode\nobreak\ \text{if }7000<k\leq N,\end{cases}\leavevmode\nobreak\ \alpha_{\zeta_{5}}(k)=\begin{cases}\displaystyle 0.006,\leavevmode\nobreak\ \text{if }1\leq k\leq 5000,\\ \displaystyle\frac{0.002}{{\rm linspace}_{(1,10,N)}(k)},\leavevmode\nobreak\ \text{if }5000<k\leq N,\end{cases}$
$\alpha_{\zeta_{6}}(k)$ $=\begin{cases}\displaystyle 0.006,\leavevmode\nobreak\ \text{if }1\leq k\leq 5000,\\ \displaystyle\frac{0.002}{{\rm linspace}_{(1,10,N)}(k)},\leavevmode\nobreak\ \text{if }5000<k\leq N,\end{cases}$
@@
Table 1: Parameters used in the simulator.

### v4

@@ 3 Continuous-time q-Function and Martingale Characterization under Tsallis Entropy
Suppose (A1), (A2) and (A3) hold. Given $(t,x)\in[0,T]\times\mathds{R}^{n}$ and $\pi,\pi^{\prime}\in\Pi_{t}$, if it holds that $F(t,x;\pi^{\prime},\pi)\geq F(t,x;\pi,\pi)$, then $J(t,x;\pi^{\prime})\geq J(t,x;\pi)$.
Moreover, in order to employ the q-learning method based on Theorem 3.3, the policy function $\pi^{\chi}$ should satisfy $\pi^{\chi}\in{\cal P}(U)$ and the consistency condition (3.6). Here, we relax these constraints and consider the following maximization problem, for $w_{1},w_{2}\geq 0$
@@
1:Initialize $\theta$, $\zeta$ and $i=1$.
@@ 5.1 The optimal portfolio liquidation problem
Consider an optimal portfolio liquidation problem in which a large investor has access both to a classical exchange and to a dark pool with adverse selection. As in Kratz and Schöneborn (2014, 2015), the trading and price formation are described as the classical exchange as a linear price impact model. The trade execution can be enforced by selling aggressively; however, this results in quadratic execution costs due to a stronger price impact. The order execution in the dark pool is modeled by a Poisson process ${N}=(N_{t})_{t\geq 0}$ with intensity parameter $\lambda>0$, where orders submitted to the dark pool are executed at the jump times of Poisson processes. The split of orders between the dark pool and the exchange is thus driven by the trade-off between execution uncertainty and price impact costs. Next, we formulate the optimal portfolio liquidation problem in detail. Consider an investor who has to liquidate an asset position $x\in\mathds{R}$ within a finite trading horizon $[0,T]$. The investor can control her trading intensity $\xi=(\xi_{t})_{t\in[0,T]}$, and can place orders $\eta=(\eta_{t})_{t\in[0,T]}$ in the dark pool. Given a trading strategy $u=(\xi_{t},\eta_{t})_{t\in[0,T]}$ (as r.c.l.l. $\mathbb{F}$-predictable processes) taking values in the policy space $U=\mathds{R}^{2}$, the asset holdings of the investor at time $t\in[0,T)$ is given by
@@
$\begin{cases}\displaystyle w(t,x):=\sup_{\pi\in\Pi}\mathbb{E}^{\mathbb{P}}\left[\int_{t}^{T}\int_{U}\left(-\kappa|u_{1}|^{2}-c|X_{s}^{u}|^{2}+\gamma l_{p}(\pi_{s}(u))\right)\pi_{s}(u)duds-g_{T}(X_{T}^{u})\Big|X_{t}=x\right],\\[11.99998pt] \displaystyle\penalty 10000\ {\rm s.t.}\penalty 10000\ X_{t}^{\pi}=x-\int_{0}^{t}\int_{\mathds{R}^{2}}u_{1}\pi_{s}(du)ds-\int_{0}^{t}\int_{\mathds{R}^{2}}u_{2}\mathcal{N}(ds,du),\quad\forall t\in[0,T].\end{cases}$ (5.5)
@@
$\begin{cases}\displaystyle V^{(\ell)}(t,x):=\sup_{\pi\in\Pi}J^{(\ell)}(t,x;\pi)\\[8.00003pt] \displaystyle\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ =\sup_{\pi\in\Pi}\mathbb{E}^{\mathbb{P}}\left[\int_{t}^{T}\int_{U}\left(-\kappa|u_{1}|^{2}-c|X_{s}^{u}|^{2}+\gamma l_{p}(\pi_{s}(u))\right)\pi_{s}(u)duds-\ell X_{T}^{2}\Big|X_{t}=x\right],\\[11.99998pt] \displaystyle\penalty 10000\ {\rm s.t.}\penalty 10000\ X_{t}^{\pi}=x-\int_{0}^{t}\int_{\mathds{R}^{2}}u_{1}\pi_{s}(du)ds-\int_{0}^{t}\int_{\mathds{R}^{2}}u_{2}\mathcal{N}(ds,du),\quad\forall t\in[0,T].\displaystyle\end{cases}$ (5.6)
@@
$\widehat{\pi}^{(\ell)}(u|t,x)=\begin{cases}\displaystyle\left(\frac{p-1}{\gamma p}\right)^{\frac{1}{p-1}}\left(\psi(t,x)-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}+\frac{\gamma}{p-1}\right)_{+}^{\frac{1}{p-1}},\penalty 10000\ p>1,\\ \\ \displaystyle\frac{\exp\left(-u_{1}V_{x}^{\ell}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}-\gamma\right)}{\int_{\mathds{R}^{2}}\exp\left(-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}-\gamma\right)du},\penalty 10000\ p=1.\end{cases}$
@@
$\widehat{\pi}^{(\ell)}(u|t,x)=\begin{cases}\displaystyle\left(\frac{p-1}{\gamma p}\right)^{\frac{1}{p-1}}\left(\psi(t,x)-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}+\frac{\gamma}{p-1}\right)_{+}^{\frac{1}{p-1}},\penalty 10000\ p>1,\\[6.80011pt] \displaystyle\frac{\exp\left(-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}-\gamma\right)}{\int_{\mathds{R}^{2}}\exp\left(-u_{1}V_{x}^{(\ell)}(t,x)+\lambda(V^{(\ell)}(t,x-u_{2})-V^{(\ell)}(t,x))-\kappa u_{1}^{2}-cx^{2}-\gamma\right)du},\penalty 10000\ p=1\end{cases}$
@@
$=\begin{cases}-\frac{\alpha^{(\ell)}(t)x}{2\kappa},\penalty 10000\ m=1,\\ \ \\ \frac{\tilde{\psi}(t)(p-1)}{2\kappa(2p-1)}+\frac{(\alpha^{(\ell)}(t)x)^{2}}{4\kappa^{2}},\penalty 10000\ m=2,\end{cases}$
@@
$\penalty 10000\ \int_{\mathds{R}^{2}}\frac{1}{p-1}\left(\pi^{(\ell)}(u|t,x)-\pi^{(\ell)}(u|t,x)^{p}\right)du$
@@
$\begin{cases}\displaystyle{\alpha^{(\ell)}_{t}(t)}=-\frac{(\alpha^{(\ell)}(t))^{2}}{2\kappa}+{\lambda}\alpha^{(\ell)}(t)+2c,\penalty 10000\ \penalty 10000\ \alpha^{(\ell)}(T)=-\ell,\\[8.00003pt] \displaystyle\beta^{(\ell)}_{t}(t)=\frac{\tilde{\psi}(t)(p-1)}{2p-1}-\gamma\left(\frac{1}{p-1}-\frac{\tilde{\psi}(t)}{(2p-1)\gamma}\right),\penalty 10000\ \penalty 10000\ \beta^{(\ell)}(T)=0.\end{cases}$
@@
For the case with $p=1$, the optimal policy $\widehat{\pi}$ given by (5.8) is a two-dimensional Gaussian distribution; while for $p>1$, the optimal policy becomes a two-dimensional $p$-Gaussian distribution with a compact support set, see Figure 1 for the illustration. The mean and variance of the policy $(u_{1},u_{2})$ are given by
${\rm mean}(u_{1},u_{2})=\left(\int_{\mathds{R}^{2}}u_{1}\widehat{\pi}(u)du,\int_{\mathds{R}^{2}}u_{2}\widehat{\pi}(u)du\right)=\left(-\frac{\alpha^{(\ell)}(t)x}{2\kappa},x\right),$ (5.9)
${\rm Var}(u_{1},u_{2})=\left(\int_{\mathds{R}^{2}}u_{1}^{2}\widehat{\pi}(u)du-\left(\int_{\mathds{R}^{2}}u_{1}\widehat{\pi}(u)du\right)^{2},\int_{\mathds{R}^{2}}u_{2}^{2}\widehat{\pi}(u)du-\left(\int_{\mathds{R}^{2}}u_{2}\widehat{\pi}(u)du\right)^{2}\right)$
$\qquad\qquad=\left(\frac{\tilde{\psi}(t)(p-1)}{2\kappa(2p-1)},-\frac{\tilde{\psi}(t)(p-1)}{\alpha^{(\ell)}(t)\lambda(2p-1)}\right).$ (5.10)
In fact, for $p>1$ and $(t,x)\in[0,T]\times\mathds{R}_{+}$, we have
@@
Figure 1: (a) The optimal policy $(u_{1},u_{2})\to\hat{\pi}(u_{1},u_{2})$ with $p=1$. (b): The optimal policy $(u_{1},u_{2})\to\hat{\pi}(u_{1},u_{2})$ with $p=2$. (c): The optimal policy $(u_{1},u_{2})\to\hat{\pi}(u_{1},u_{2})$ with $p=3$. The model parameters are set to be $\lambda=1,\penalty 10000\ \kappa=1,\penalty 10000\ c=1,\penalty 10000\ \gamma=1,\penalty 10000\ t=1,\penalty 10000\ T=2,\penalty 10000\ x=5$.
@@
which can also be observed from Figure 2 and Figure 3. It demonstrates that the entropy index $p$ can essentially determine the compact support of the exploratory policy. As $p\to 1$ this framework recovers the Shannon entropy limit. Therefore, while a Tsallis-regularized algorithm may not offer a direct advantage in convergence or sample efficiency over its Shannon counterpart, it brings a practical benefit of bounded sampling of actions from the compact support, which might be crucial in some applications.
Figure 2: (a): The compact support set of $u_{1}$. (b) The compact support set of $u_{2}$. The model parameters are set to be $\lambda=1,\penalty 10000\ \kappa=1,\penalty 10000\ c=1,\penalty 10000\ \gamma=0.001,\penalty 10000\ t=1,\penalty 10000\ T=2,\penalty 10000\ x=5$.
Figure 3: (a): The compact support set of $u_{1}$. (b) The compact support set of $u_{2}$. The model parameters are set to be $\lambda=1,\penalty 10000\ \kappa=1,\penalty 10000\ c=1,\penalty 10000\ \gamma=0.1,\penalty 10000\ t=1,\penalty 10000\ T=2,\penalty 10000\ x=5$.
@@
which yields the convergence of the optimal trading policy to a deterministic strategy that $(\xi_{t},\eta_{t})\xrightarrow[\gamma\to 0]{L^{2}}\left(\frac{\alpha^{*}(t)x}{2\kappa},x\right)$ for all $(t,x)\in[0,T]\times\mathds{R}$. Note that $(\frac{\alpha^{*}(t)x}{2\kappa},x)$ is the optimal strategy to the classical stochastic control problem ($\gamma=0$). That is, the exploratory problem reduces to the classical problem as temperature parameter $\gamma\to 0$.
@@
Simulator: In what follows, we apply Algorithm 1 with the above parameterized value function and q-function. To generate sample trajectories, we first apply the acceptance-rejection sampling method (Flury 1990) to generate the control pair $(u^{1}_{t},u^{2}_{t})$ from the $p$-Gaussian distribution with density function given by (5.14) at time $t$. Then, the control pair $(u^{1}_{t},u^{2}_{t})$ is used to the following simulator
@@
$\alpha_{\theta_{1}}(k)$ $=\begin{cases}\displaystyle 0.01,\penalty 10000\ \text{if }1\leq k\leq 2500,\\ \displaystyle\frac{0.001}{{\rm linspace}_{(1,20,N)}(k)},\penalty 10000\ \text{if }2500<k\leq N,\end{cases}\penalty 10000\ \alpha_{\theta_{2}}(k)=\begin{cases}\displaystyle 0.005,\penalty 10000\ \text{if }1\leq k\leq 4000,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,100,N)}(k)},\penalty 10000\ \text{if }4000<k\leq N,\end{cases}$
$\alpha_{\theta_{3}}(k)$ $=\begin{cases}\displaystyle 0.01,\penalty 10000\ \text{if }1\leq k\leq 4000,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,20,N)}(k)},\penalty 10000\ \text{if }4000<k\leq N,\end{cases}\penalty 10000\ \alpha_{\theta_{4}}(k)=\begin{cases}\displaystyle 0.03,\penalty 10000\ \text{if }1\leq k\leq 3000,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,20,N)}(k)},\penalty 10000\ \text{if }3000<k\leq N,\end{cases}$
$\alpha_{\theta_{5}}(k)$ $=\begin{cases}\displaystyle 0.05,\penalty 10000\ \text{if }1\leq k\leq 3000,\\ \displaystyle\frac{0.0005}{{\rm linspace}_{(1,20,N)}(k)},\penalty 10000\ \text{if }3000<k\leq N,\end{cases}\penalty 10000\ \alpha_{\zeta_{1}}(k)=\begin{cases}\displaystyle 0.03,\penalty 10000\ \text{if }1\leq k\leq 3500,\\ \displaystyle\frac{0.00135}{{\rm linspace}_{(1,10,N)}(k)},\penalty 10000\ \text{if }3500<k\leq N,\end{cases}$
$\alpha_{\zeta_{2}}(k)$ $=\begin{cases}\displaystyle 0.1,\penalty 10000\ \text{if }1\leq k\leq 3500,\\ \displaystyle\frac{0.0002}{{\rm linspace}_{(1,500,N)}(k)},\penalty 10000\ \text{if }3500<k\leq N,\end{cases}\penalty 10000\ \alpha_{\zeta_{3}}(k)=\begin{cases}\displaystyle 0.1,\penalty 10000\ \text{if }1\leq k\leq 2000,\\ \displaystyle 0.002,\penalty 10000\ \text{if }2000<k\geq 5000,\\ \displaystyle\frac{0.0005}{{\rm linspace}_{(1,20,N)}(k)},\penalty 10000\ \text{if }k\geq 5000,\end{cases}$
$\alpha_{\zeta_{4}}(k)$ $=\begin{cases}\displaystyle 0.005,\penalty 10000\ \text{if }1\leq k\leq 7000,\\ \displaystyle\frac{0.001}{{\rm linspace}_{(1,100,N)}(k)},\penalty 10000\ \text{if }7000<k\leq N,\end{cases}\penalty 10000\ \alpha_{\zeta_{5}}(k)=\begin{cases}\displaystyle 0.006,\penalty 10000\ \text{if }1\leq k\leq 5000,\\ \displaystyle\frac{0.002}{{\rm linspace}_{(1,10,N)}(k)},\penalty 10000\ \text{if }5000<k\leq N,\end{cases}$
$\alpha_{\zeta_{6}}(k)$ $=\begin{cases}\displaystyle 0.006,\penalty 10000\ \text{if }1\leq k\leq 5000,\\ \displaystyle\frac{0.002}{{\rm linspace}_{(1,10,N)}(k)},\penalty 10000\ \text{if }5000<k\leq N,\end{cases}$
@@
Table 1: Parameters